OpenStudy (anonymous):

cos theta-tan theta cos theta=0 a. 0, pi/4, pi, 5pi/4 b.pi/4, 5pi/4 c. pi/2, 3pi/4, 3pi/2, 7pi/4 d. pi/2, 7pi/6, 3pi/2, 11pi/6 I believe the answer is b but im not positive

OpenStudy (tanya123):

B.pi/4, 5pi/4

OpenStudy (anonymous):

are you positive? @tanya123

OpenStudy (tanya123):

Yes.

OpenStudy (anonymous):

could you help me with another? if you dont mind? @tanya123

OpenStudy (anonymous):

4cos theta + 1= 2cos theta

OpenStudy (tanya123):

Sorry I don't know this one, but I think he could help you @jim_thompson5910

OpenStudy (anonymous):

@jim_thompson5910 can you help?

jimthompson5910 (jim_thompson5910):

think of cos(theta) as 'z' so we really have this equation 4z + 1 = 2z solve for z to get ???

OpenStudy (anonymous):

2?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

4z + 1 = 2z 4*2 + 1 = 2*2 8 + 1 = 4 9 = 4 ... that equation is false, so z = 2 is not the solution

OpenStudy (anonymous):

oh! ohkay

OpenStudy (anonymous):

wait, so would that mean there is no answer?

OpenStudy (anonymous):

because thats not one of my choices?

jimthompson5910 (jim_thompson5910):

4z + 1 = 2z 4z = 2z - 1 4z - 2z = -1 ... ... ... z = ??

jimthompson5910 (jim_thompson5910):

don't worry about the final answer right now, we're just focusing on solving for z

OpenStudy (anonymous):

you know what, im missing part of this problem. what values for theta(0greater than or equal to theta greater than or equal to 2 pi) satisfy the equation im not sure if that makes a difference

jimthompson5910 (jim_thompson5910):

it does, but not for this bit when solving for z

OpenStudy (anonymous):

2z=-1 z=-2

jimthompson5910 (jim_thompson5910):

2z = -1 is right z = -2 is NOT correct

OpenStudy (anonymous):

hmm -1/2?

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

we made z = cos(theta) so what we really have is cos(theta) = -1/2

jimthompson5910 (jim_thompson5910):

use a unit circle to solve for theta in cos(theta) = -1/2

OpenStudy (anonymous):

im sorry, I don't really understand how use the unit circle

jimthompson5910 (jim_thompson5910):

do you have a unit circle with you?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

look for points on the unit circle that have an x coordinate of -1/2

jimthompson5910 (jim_thompson5910):

what's the angle that's associated with them (in radians)?

OpenStudy (anonymous):

in quadrant III and II

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

you're in the right quadrants

OpenStudy (anonymous):

awesome, so what next

jimthompson5910 (jim_thompson5910):

locate the point in quadrant 2 has an x coordinate of -1/2 what is the angle of this point?

OpenStudy (anonymous):

120 and 240

jimthompson5910 (jim_thompson5910):

120 is in quadrant 2 240 in quadrant 3 those are the degree angles what about the radian angles?

OpenStudy (anonymous):

2pi/3 and 4pi/3

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

so is that my answer?

jimthompson5910 (jim_thompson5910):

yes, those are the solutions to 4cos(theta) + 1 = 2cos(theta)

jimthompson5910 (jim_thompson5910):

where 0 <= theta < 2pi

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

could you check the answer to my first question to make sure it is right?

jimthompson5910 (jim_thompson5910):

can you post it again (with the answer you got)

OpenStudy (anonymous):

yes of course